OpenStudy (anonymous):

x=log 5 base 3 y=log25 base 16 find relation between x and y....

OpenStudy (anonymous):

3^x=5 -----> 5 x 3^x = 25 16^y=25 5x3^x=16^y alternately, 3^x=5 ----> 3^2x=25 3^2x=16^y take log on both sides 2xlog3 = ylog16

OpenStudy (anonymous):

hmm mark it as answered please :)